### v2

# RSPO: Regularized Self-Play Alignment of Large Language Models

Abstract
Self-play alignment has emerged as an effective approach for fine-tuning large language models (LLMs), formulating preference optimization as a two-player game. However, the regularization with respect to the reference policy, which is crucial for mitigating over-optimization, has been insufficiently investigated in self-play alignment. To study the impact of different regularization strategies, we propose Regularized Self-Play Policy Optimization (RSPO), a general and modular framework that unifies prior methods and enables simple plug-and-play integration of various regularizers, meanwhile preserving convergence to Nash equilibrium of the corresponding regularized game.Our empirical study involving over $120$ fine-tuned Mistral-7B-Instruct models reveals that forward KL divergence regularization reduces response length, whereas reverse KL divergence markedly improves raw win rates. Crucially, RSPO regularized with a linear combination of forward and reverse KL divergence significantly boosts the length-controlled win rate on AlpacaEval-2 from $28.5\%$ (unregularized self-play, SPPO) to $35.4\%$, and consistently demonstrates superior performance on Arena-Hard, MT-Bench, ArmoRM scores, and response diversity. Combining simplicity, convergence guarantees, and significant empirical gains, RSPO offers a strong foundation for exploring regularized self-play in language model alignment.
1 Introduction
Self-play is a line of work conducting iterative self-competition of models, which has been demonstrated as an effective approach for improving AI systems (Goodfellow et al., 2020; Wang et al., 2022), particularly in strategic decision-making problems (Silver et al., 2016; Heinrich and Silver, 2016; Pinto et al., 2017; Brown and Sandholm, 2018). In the human alignment of LLMs, self-play recently started to be used and has shown superior empirical performance than other iterative Reinforcement Learning from Human Feedback (RLHF) methods on popular benchmarks (Dubois et al., 2024; Jiang et al., 2024; Wu et al., 2024; Rosset et al., 2024). By formulating the preference optimization problem as a two-player game, self-play alignment methods seek to identify a Nash Equilibrium (NE) of the game in which utility is determined by a general preference model (Azar et al., 2024; Munos et al., 2023; Calandriello et al., 2024). This NE is regarded as the most aligned LLM policy achieved without Bradley-Terry (BT) reward modeling (David, 1963), which has shown under-performance compared to general preference modeling (Ye et al., 2024).
Figure 1: RSPO is flexible for general regularization. The estimation of Nash-MD policy update direction $\partial_{\pi}\mathbb{P}(\pi\succ\pi_{t}^{\mu})$ requires samples from geometric mixture policy $\pi_{t}^{\mu}$. Such update approach is only compatible with reverse KL divergence for regularization.
Despite the significant empirical improvements achieved through self-play, the impact of regularization to the reference policy—commonly used in RLHF to mitigate over-optimization—has received insufficient investigation in self-play alignment. Most existing self-play methods completely lack explicit regularization (Wu et al., 2024; Rosset et al., 2024; Swamy et al., 2024; Wang et al., 2024a; Gao et al., 2024). In practice, unregularized self-play is also susceptible to over-optimization, particularly when the preference model is inaccurate or misspecified. Although a few recent self-play approaches like Nash-MD (Munos et al., 2023) incorporate reverse KL divergence as a regularization penalty (Calandriello et al., 2024; Wang et al., 2024a; Zhang et al., 2024), it remains unclear whether reverse KL is optimal for alignment, and the broader impact of alternative regularization strategies in self-play remains insufficiently explored. Moreover, the extension of current approaches to general forms of regularization is challenging, as their training protocols are intrinsically reliant on the reverse KL divergence for regularization Munos et al. (2023) (see Figure 1).
In this work, we introduce a flexible framework for incorporating diverse regularization methods into self-play alignment, termed Regularized Self-Play Policy Optimization (RSPO):
  • RSPO offers a simple way to apply various regularization strategies in self-play by directly adding the regularization term to our proposed unified self-play loss function, while maintaining last-iterate convergence to NE of the corresponding regularized preference optimization game.
  • We conduct empirical analysis on over $120$ fine-tuned Mistral-7B-Instrct models, which reveals distinct effects of different regularizations: forward KL regularization reduces the response length, whereas reverse KL regularization significantly improves the raw win rate.
  • Consequently, we propose a linear combination of forward and reverse KL divergence, yielding a substantial improvement over the unregularized self-play alignment method (SPPO (Wu et al., 2024)) on popular benchmarks: AlpacaEval-2.0, Arena-Hard-v0.1, and MT-bench. Particularly on AlpacaEval-2.0, RSPO outperforms SPPO with improvement of $6.9\%$ in length-controlled win rate (LCWR).
  • Finally, we provide quantitative evaluations on other aspects of the generated responses, including diversity via self-BLEU (Zhu et al., 2018), and instruction following, truthfulness, honesty, helpfulness via ArmoRM (Wang et al., 2024b). Regularization demonstrates improvements on almost all aspects of the responses, indicating the comprehensive effectiveness of regularization in self-play alignment.
2 Preliminaries
We denote a prompt as $x$, a response as $y$, and a LLM policy as $\pi(y|x)$, where $\pi(\cdot|x)\in\Delta_{\mathcal{Y}}$. We denote the set of all prompts as $\mathcal{X}$, and the set of all responses as $\mathcal{Y}=\{y^{0},y^{1},\cdots\}$. We use $\Delta_{\mathcal{Y}}$ to denote the probability simplex over the responses given a specific prompt. We parametrize the LLM policy $\pi$ as $\pi_{\theta}$. The reference policy is an LLM denoted as $\mu\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$. For notational brevity, we remove the dependence of policy $\pi$ and loss functions on the prompt $x$ throughout the paper.
2.1 Game-Theoretic Preference Optimization
We study the preference optimization problem in an online setting by formulating it as a two-player max-min game, as studied in previous self-play works (Wu et al., 2024). The players are two LLMs whose strategies are LLM policies, denoted as max-player $\pi$ and min-player $\pi^{\prime}$. The utility of the max-player is expressed as the preference of itself over the min-player:
[…]
where $u:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow\mathbb{R}$ is linear in $\pi$ and $\pi^{\prime}$; $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$ is a general preference model that quantifies the preference of $y$ over $y^{\prime}$ given a prompt. We extend the notation $\mathbb{P}(y\succ\pi^{\prime})=\mathbb{E}_{y^{\prime}\sim\pi^{\prime}}[\mathbb{P}(y\succ y^{\prime})]$. The objective is finding a NE policy $\pi^{*}$ of the preference model:
$(\pi^{*},\pi^{*})=\arg\max_{\pi}\min_{\pi^{\prime}}\mathbb{P}(\pi\succ\pi^{\prime}).$ (2)
Therefore, an NE strategy $\pi^{*}$ is an LLM that can generate the most preferred responses in expectation, thus achieving human alignment based on the preference model. Most existing self-play alignment methods aim to solve this NE following Algorithm 1 (Wu et al., 2024; Rosset et al., 2024; Swamy et al., 2024; Wang et al., 2024a).
2.2 Preference Optimization via Multiplicative Weights Update
An effective self-play method to solve the preference optimization game in Equation 2 is Self-Play Policy Optimization (SPPO) (Wu et al., 2024). SPPO derives its loss function from the iterative no-regret learning algorithm, Multiplicative Weights Update (MWU) (Freund and Schapire, 1997). Specifically in a game setting, denote learning rate as $\eta$, and normalization constant $Z(\pi_{t})$. In any iteration $t$, the policy update $\forall y\in\mathcal{Y}$ is $\pi_{t+1}(y)=\pi_{t}(y)\cdot{\exp\big{(}\eta\mathbb{E}_{y^{\prime}\sim\pi_{t}}[u({y};y^{\prime})]\big{)}}/{Z(\pi_{t})},$ where $u(y;y^{\prime})$ is the utility function defined in Equation (1), with $y$ treated as a pure strategy.
The practical loss function of SPPO for policy update is then derived according to MWU:
$\mathcal{L}_{\text{SPPO}}(\theta)=\mathbb{E}_{y\sim\pi_{t}}\Big{[}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\Big{(}\eta\mathbb{P}({y}\succ\pi_{t})-\log Z(\pi_{t})\Big{)}\Big{]}^{2}.$ (3)
SPPO converges to the NE of the preference optimization game in Equation 2. However, after running multiple iterations, the deviation of the policy $\pi_{\theta}$ from $\mu$ can be large. Such deviation is particularly problematic when the preference model is only accurate at evaluating responses sampled from the reference policy (Munos et al., 2023). Furthermore, in aligning LLMs in practice, the preference model is typically a surrogate $\hat{\mathbb{P}}$, such as PairRM (Jiang et al., 2023a), which may be misspecified at some out-of-distribution responses and inaccurate due to estimation error or limited model expressiveness (PairRM is only a 0.4B model), causing over-optimization problem. Regularizing the policy optimization to a reference SFT model, which is typically trained on high-quality data (Ouyang et al., 2022), can mitigate the problem. We provide a synthetic example in Appendix D.1 to demonstrate this problem.
2.3 Regularized Preference Optimization Game with Reference Policy
To address the regularization in self-play, we adopt the objective in Nash Learning from Human Feedback (Munos et al., 2023), and extend the KL divergence regularization to a general regularization function, to penalize the deviation from the reference policy. We define a convex regularization function $R:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$, where $R(\pi,\mu)$ measures the distance between $\pi$ and the reference model $\mu$, such as KL divergence $D_{\text{KL}}(\pi\|\mu)$. Denote regularization temperature as $\tau$, the objective becomes to optimize a regularized preference model by solving the NE $(\pi^{*},\pi^{*})$ of the regularized game, where the utility of max player is still $u(\pi;\pi^{\prime})=\mathbb{P}(\pi\succ\pi^{\prime})$:
$\arg\max_{\pi}\min_{\pi^{\prime}}\mathbb{P}(\pi\succ\pi^{\prime})-\tau R(\pi,\mu)+\tau R(\pi^{\prime},\mu).$ (4)
We provide proof of the existence and uniqueness of this NE in Appendix A.1. A few recent methods leverage Mirror Descent (MD), which is also in a self-play manner, to find a regularized NE in Equation 4 with last-iterate policy (Munos et al., 2023; Calandriello et al., 2024; Zhang et al., 2024).
However, these MD-based methods are only compatible with the reverse KL divergence regularizer, and are non-trivial to extend to general divergence. For instance, Nash-MD [^1] Throughout the paper, regularization specifically refers to the deviation of $\pi$ from $\mu$, rather than from $\pi_{t}$. addresses the reverse KL regularization of $\pi$ and $\mu$ requiring responses generated from a geometric mixture policy $\pi_{t}^{\mu}(y)\propto{\pi_{t}(y)^{1-\eta\tau}\mu(y)^{\eta\tau}}$ (Munos et al., 2023), which is inherently compatible only with reverse KL divergence:
$\pi_{t+1}=\arg\min_{\pi}-\eta\mathbb{E}_{\pi}[\nabla_{\pi}u(\pi_{t};{\color[rgb]{1,0,0}\pi^{\mu}_{t}})]+D_{\text{KL}}(\pi\|\pi_{t}^{\mu}).$ (5)
Therefore, while the LLMs optimized via existing self-play methods exhibit empirical improvement, they all have limited regularization of $\pi$ and $\mu$. The potential benefits of alternative regularization, such as adopting other $f$-divergences than reverse KL, remain unexplored.
3 Regularized Self-Play Policy Optimization
We propose a framework for regularized self-play alignment, namely Regularized Self-Play Policy Optimization (RSPO). RSPO is simple and flexible for regularization, and provably convergent to Nash Equilibrium. The loss function of RSPO $\mathcal{L}_{\text{RSPO}}$ is defined as the sum of a mean-squared self-play loss and a weighted regularization term:
[…]
where $G({y},\pi_{t},\mu)$, $B(\pi_{t},\mu)$, and $R(\pi_{\theta},\mu)$ are configurable components. First, $G:\mathcal{Y}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$ defines the update direction of $\pi_{\theta}$, which can be set as the gradient of a utility function to guide the policy update towards increasing the utility. Second, the baseline function $B:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$ is for variance-reduction of $G$, similar to the baseline in REINFORCE (Williams, 1992). Lastly, $R:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow\mathbb{R}$ is the regularization function. The coefficient $\lambda$ is the regularization temperature. The first Mean Square Error term in Equation 6 can be interpreted as a self-play loss of conducting exponentiated gradient descent (Beck and Teboulle, 2003).
RSPO is a modular framework offering a simple way to introduce regularization into self-play alignment with only an additional term in the loss. RSPO offers the simplicity and flexibility to incorporate various regularization methods into self-play-based preference optimization methods. Additionally, we show in Section 3.1 that RSPO can generalize existing unregularized self-play methods without external regularization $R$. Thus, regularizing existing methods requires no change to their original loss functions or hyperparameters, but simply adding an external plug-and-play regularization to their loss function and tuning the temperature $\lambda$.
In practice, we set baseline function $B=\tfrac{1}{2}$ following Nash-MD and SPPO, and the update direction $G$ to be the gradient of the preference against $\pi_{t}$, $\forall y\in\mathcal{Y}$:
$G({y},\pi_{t},\mu)=\partial_{\pi(y)}\mathbb{P}(\pi\succ\pi_{t})=\mathbb{P}(y\succ\pi_{t}).$ (7)
We execute Algorithm 1 by applying the following RSPO loss with any regularization $R$ of interests:
$\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}),B=\tfrac{1}{2},R\big{)}.$ (8)
In theory, $B$ helps minimize the variance of $G$ the most when $B=\mathbb{E}_{y\sim\pi_{t}}[G(y,\pi_{t},\mu)]$. But in preference optimization, due to the typically small minibatch size, the estimation error of the mean of $G$ could be large, leading to additional estimation error of the loss. Thus, we also set the baseline value for variance reduction to be a constant $\tfrac{1}{2}$, the mean value of $G$ when the algorithm converged. For the implementation of various divergence-based regularization, refer to Appendix C.3.
In the following sections, we first illustrate the generalizable formulation of RSPO, so that it can be implemented without modifying the existing self-play component. We then establish theoretical convergence guarantees for RSPO grounded in Mirror Descent theory.
3.1 Generalizing Existing Self-Play Methods
In this section, we show how RSPO generalize existing self-play methods, which showcase (1) implementing RSPO requires only one additional term to existing self-play loss functions; (2) the limitation of existing regularized methods. First, the unregularized self-play method SPPO (Wu et al., 2024) has a loss function defined in Equation 3 equivalent to RSPO without external regularization:
[…]
According to Equation 8 and Equation 9, $\mathcal{L}_{\text{RSPO}}=\mathcal{L}_{\text{SPPO}}+\lambda R(\pi_{\theta},\mu)$, i.e. the implementation of RSPO is equivalent to directly add the regularization $R$ to the loss function of SPPO (Equation 3). This implies that the additional regularization term becomes plug-and-play, requiring minimal changes to existing training pipeline.
In addition, existing regularized methods can be generalized by $\mathcal{L}_{\text{RSPO}}$ (derivations in Appendix A.2):
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}(\theta)=$ $\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu}),B=\frac{1}{2},R=D_{\text{KL}}(\pi_{\theta}\|\mu)\big{)}$ (10)
$=$ $\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\Big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu})-\tau\log\frac{\pi_{t}(y)}{\mu(y)},B=\frac{1}{2},R=0\Big{)}.$ (11)
Equation 10 verifies our summarization shown in Figure 1. The convergence guarantee of Nash-MD (Munos et al., 2023, Lemma 2) requires the policy updated with Equation 5, which is specifically designed for reverse KL regularization, as other $R$ can not be merged with $D_{\text{KL}}(\pi\|\mu)$ to a regularization w.r.t. geometric mixture $\pi^{\mu}_{t}$. Additionally, Equation 11 demonstrates that RSPO enables to even add extra regularization to existing regularized self-play methods.
3.2 Theoretical Guarantees
In this section, we examine the theoretical properties of RSPO, with a particular emphasis on its convergence guarantee. We adopt Mirror Descent (MD) as the foundational framework, given its well-established last-iterate convergence to the NE.
We build upon Magnetic Mirror Descent (MMD) (Sokota et al., 2022), a specialized variant of MD that guarantees convergence to a reverse-KL-regularized NE. To generalize beyond reverse-KL regularization, we introduce Generalized Magnetic Mirror Descent (GMMD), which can accommodate a broader class of regularization techniques. By demonstrating that optimizing the RSPO loss is equivalent to performing reinforcement learning (RL) within the GMMD framework, we establish a formal connection between RSPO and GMMD. This connection ensures the last-iterate convergence of RSPO to the NE of the corresponding regularized game.
Tabular GMMD. Denote the utility function of the game as $U$, define $G$ as the element of the vector of partial derivatives of $U$ w.r.t. policy:
[…]
Then in iteration $t$, GMMD updates policy as
$\pi_{t+1}=\arg\min_{\pi}-\eta\mathbb{E}_{\pi}[G(y;\pi_{t})]+B_{\psi}(\pi;\pi_{t})+\tau R(\pi,\mu),$ (13)
where $\tau$ is regularization temperature, $R$ is a general regularization function, serving as a “magnet” to attract $\pi$ to $\mu$ during policy updating. $B_{\psi}$ is the Bregman Divergence generated by a convex potential function $\psi$ (Bregman, 1967).
Notably, the vanilla Magnetic Mirror Descent limits $R$ to be the same regularization method of $\pi$ and $\pi_{t}$, i.e., $R=B_{\psi}$ (Sokota et al., 2022, Section 3.2); whereas in this paper we aim at a general regularizer of $\pi$ and $\mu$, which could be different from $B_{\psi}$, and study the effects of different regularization methods.
Proposition 3.1 (Last-iterate Convergence).
If $R(\cdot,\mu)$ is $1$-strongly convex relative to $\psi$, $\eta\leq\tau$, and $U$ is linear, then policy updated by GMMD in Equation 13 has last-iterate convergence to the following regularized NE $\max_{\pi}\min_{\pi^{\prime}}U(\pi;\pi^{\prime})-\tau R(\pi,\mu)+\tau R(\pi^{\prime},\mu).$
Proposition 3.1 is a direct application of Theorem 3.4 by Sokota et al. (2022), which guarantees the last-iterate convergence of GMMD to the NE of a regularized game (Proof in Appendix A.3).
Deep RL Implementation of GMMD. To adapt GMMD to preference optimization problems, RL techniques are commonly employed as practical implementations, as for many MD update (Tomar et al., 2020; Munos et al., 2023; Wang et al., 2024a). Define the loss function of conducting GMMD in preference optimization as
[…]
Here, we set the Bregman divergence to Reverse KL in preference optimization as in previous works (Munos et al., 2023; Zhang et al., 2024). The gradient estimation of $\mathcal{L}_{\text{GMMD}}(\theta)$ for policy updates is required since the expectation in the first term is dependent on $\pi_{\theta}$. Following Policy Gradient (PG) theorem (Sutton et al., 1999), the PG of GMMD is equal to $\nabla_{\theta}\mathcal{L}_{\text{RSPO}}(\theta)$ up to multiplying a constant:
$\nabla_{\theta}\mathcal{L}_{\text{GMMD}}(\theta)=\mathbb{E}_{y\sim\pi_{\theta}}\bigg{[}\nabla_{\theta}\log\pi_{\theta}(y)\bigg{(}-\eta G(y;\pi_{t})+\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}+B\bigg{)}\bigg{]}+\tau\nabla_{\theta}R(\pi_{\theta},\mu),$ (15)
where $B$ is a baseline function to reduce the variance as in REINFORCE (Williams, 1992). We set $B$ independent to $\theta$ so that adding $B$ does not affect the value of Equation 14, due to $\mathbb{E}_{y\sim\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(y)\cdot\eta B]=\eta B\nabla_{\theta}\mathbb{E}_{y\sim\pi_{\theta}}[1]=0$.
Due to the equivalence between RSPO and GMMD, we provide the convergence guarantee for our practical implementation of RSPO (Equation 13), to the Nash equilibrium of the regularized preference optimization game as follows (Proof in Appendix A.4).
Corollary 3.2.
Self-play following Algorithm 1 with the RSPO loss function in Equation 8 and regularizer $R$ satisfying the assumption in Proposition 3.1, has last-iterate convergence to the NE of the regularized preference optimization game, as described in Equation 4.
RSPO guarantees NE convergence while allowing flexible regularization strategies, making it a robust extension of self-play optimization. In summary, the proposed RSPO framework provides a generalized approach that simplifies the incorporation of regularization into existing self-play methods while maintaining theoretical soundness.
4 Experiments
In this section, we answer the following important questions of regularization in the self-play alignment of Large Language Models (LLMs) by testing on various popular benchmarks:
  • Q1: Does regularization improve the performance of self-play alignment (Sec. 4.1)?
  • Q2: Which regularization method is the most effective in self-play alignment (Sec. 4.2)?
  • Q3: What additional advantages can be obtained by regularization in self-play (Sec. 4.3)?
Experiment Setup. We evaluate our methods mainly on benchmarks AlpacaEval (Dubois et al., 2024), Arena-Hard (Li et al., 2024), and MT-Bench (Zheng et al., 2023), and test the response generation diversity and quality via self-BLEU (Zhu et al., 2018) and ArmoRM (Wang et al., 2024b), respectively. We follow the experiment setup of SPPO and Snorkel-Mistral-PairRM-DPO (Snorkel) (Tran et al., 2023) to examine our regularization methods, where Snorkel is based on iterative DPO and has achieved strong performance on AlpacaEval. Our reference policy model and base model are both Mistral-7B-Instruct-v0.2. Since iterative self-play methods require no response data for training, we only use the prompts of the Ultrafeedback dataset (Cui et al., 2023), whose size is $\sim 60$K. Following SPPO and Snorkel, we split the prompts into three subsets and use only one subset per iteration to prevent over-fitting. To understand the later-iterate performance of self-play, in section 4.1, we also train on the single fold of the prompts iteratively. We use a 0.4B response-pair-wise preference model PairRM (Jiang et al., 2023a), evaluated as comparable to 10×10$\times$10 × larger reward/preference models (Cui et al., 2023).
Implementations and Baselines. The implementation of self-play methods follows Algorithm 1. In each iteration, given response-pair-wise preference from PairRM and $K=5$ number of response samples from the current policy, we estimate the policies’ preference $\mathbb{P}(\pi\succ\pi_{t})$ and regularization via Monte-Carlo estimation to compute the loss function. We replicate SPPO with the default hyper-parameters and extend it to $9$ iterations. We implement RSPO as described in Corollary 3.2. The implementation of regularizations in RSPO is demonstrated in Appendix C.3 using the $K$ samples. We report some of the baseline results from the previous papers, including SPPO, Snorkel (Mistral-PairRM-DPO) (Tran et al., 2023), Mistral-7B (Instruct-v0.2) (Jiang et al., 2023b), iterative DPO by Wu et al. (2024), and SimPO Meng et al. (2024). Since the SPPO paper only provides results across $3$ iterations (Wu et al., 2024), we replicate SPPO as an important baseline to study the performance across more than $3$ iterations.
Figure 2: Left: Length-controlled win rate (LCWR) across iterations for unregularized self-play method: SPPO, SPPO trained on a subset of the data: SPPO (subset), and reverse-KL regularized RSPO: RSPO (Rev. KL). SPPO starts to degrade from iteration 3 due to overoptimization, RSPO with reverse KL regularization mitigates it. Right: LCWR of SPPO and RSPOs in AlpacaEval-2.0. RSPO (IS-For.+Rev.) outperforms SPPO and others including reverse KL regularization.
4.1 Effectiveness of Regularization
In this section, we assess the effectiveness of regularization primarily by comparing the performance of unregularized and regularized self-play methods. We first examine the over-optimization issue inherent in practical self-play alignment by extending the execution of SPPO to Iteration 5. As depicted in Figure 2 (left), a decline in performance appears during the later iterations of SPPO. We hypothesize that this behavior arises from the practical challenges associated with over-optimization.
We then present comprehensive results across three widely used benchmarks (Table 1). RSPO with forward and reverse KL regularization, consistently outperforms the unregularized baseline (SPPO) [^2] We report our replicated testing of the published SPPO Iter3 model (link) on Arena-Hard benchmark. Thus, it is different from the result presented in the original paper of SPPO (Wu et al., 2024)., and iterative DPO in iteration 3, and the strong offline method SimPO across all benchmarks, with a clear performance margin. These results underscore the importance of incorporating regularization into self-play alignment. We hypothesize that the effectiveness of regularization arises from the continued utility of the reference policy during optimization, which provides stable guidance and helps mitigate inaccuracies or misspecifications in the general preference model (PairRM).
Methods
(Base Model: Mistral-7B-Instruct)
AlpacaEval-2
[…]
Arena-Hard
Auto-v0.1
MT-Bench
Mistral-7B-Instruct (Jiang et al., 2023b) 17.1 12.6 7.51
Snorkel (Iterative DPO) (Tran et al., 2023) 26.4 20.7 7.58
SPPO Iter3 (Wu et al., 2024) 28.5 19.2 7.59
SimPO (Meng et al., 2024) 32.1 21.0 7.60
RSPO (IS-For.+Rev.) Iter3 35.4 22.9 7.75
Table 1: Comprehensive comparisons on three popular benchmarks of baselines, and our strongest model. RSPO with Importance-Sampling-based Forward KL ($\lambda_{1}=0.1$) + Reverse KL ($\lambda_{2}=0.5$) divergence as regularization outperforms baselines on all benchmarks with a clear margin.
Model AlpacaEval 2.0
LC Win Rate Win Rate Avg. Len
Mistral-7B 17.11 14.72 1676
Snorkel 26.39 30.22 2736
SimPO 32.1 34.8 2193
DPO Iter1 23.81 20.44 1723
DPO Iter2 24.23 24.46 2028
DPO Iter3 22.30 23.39 2189
SPPO Iter1 24.79 23.51 1855
SPPO Iter2 26.89 27.62 2019
SPPO Iter3 28.53 31.02 2163
[…]
RSPO Iter1 23.16 21.06 1763
RSPO Iter2 27.91 27.38 1992
RSPO Iter3 35.44 38.31 2286
Table 2: AlpacaEval LCWR of iterative methods. RSPO shows fast improvement over iterations.
In Table 2, we further contrast the performance dynamics across iterations of methods: unregularized self-play method, (SPPO), other iterative methods, and the best RSPO, namely RSPO (For.+Rev.), regularized by the linear combination of Forward KL and Reverse KL divergence with temperatures of $0.1$ and $0.5$, respectively. The comparative results reveal that regularization enhances the SPPO win rate from $31.02\%$ to $38.31\%$, and the LC win rate increases from $28.53\%$ to $35.44\%$ in iteration $3$. Notably, in the first iteration, reg. SPPO exhibits a slightly lower LC win rate, potentially attributable to the influence of strong regularization. However, subsequent iterations show a marked improvement, with the LC win rate of reg. SPPO increases by up to $7.53\%$ within a single iteration. In summary, Table 2 underscores the effectiveness of regularization in self-play optimization.
[…]
4.2 Impact of Different Regularizations
We then study the effect of applying different regularization $R$ in RSPO. To obtain a well-regularized self-play, the tuning of regularization temperature $\lambda$ is necessary. An ablation study of the regularization temperature of different methods is shown in Figure 3. According to the figure, the response length increases along with the temperature in reverse KL divergence and Chi-square divergence regularized RSPO. However, both the length and win rate are decreased with stronger regularization via Forward KL divergence, implemented using importance sampling. We attribute the decreasing win rate to the violation of relative convexity assumption (A.1), and the length reduction to the intrinsic mass-averaging property of forward-KL divergence divergence when used for regularization.
In particular, the raw win rate analysis highlights reverse KL divergence as a crucial factor in enhancing self-play performance. We attribute the observed effect to the inherent mode-seeking behavior of reverse KL divergence. Given that forward KL divergence tends to reduce response length while reverse KL divergence yields significant improvements, we adopt a linear combination of both. This approach is designed to balance their complementary effects, ultimately optimizing for a higher LCWR (RSPO (IS-For. + Rev.) in Figure 2 Right). The hyperparameters provided in Table 6.
Figure 3: Ablation study on regularization temperature $\lambda$ of RSPO conducted on AlpacaEval 2.0 prompts. We evaluate how the average response length and raw WR are affected by the regularization temperature. Higher temperature of forward KL leads to shorter response length.
4.3 Response Diversity and Other Aspects
Regularization Iteration AlpacaEval 2.0 Dataset
LCWR $\uparrow$ Self-BLEU $\downarrow$
$\times$ 1 24.79 0.751
2 26.89 0.754
3 28.53 0.758
IS-Forward KL + Reverse KL 1 23.16 0.747
2 27.91 0.743
3 35.44 0.714
Reverse KL 1 25.52 0.747
2 32.26 0.730
3 34.21 0.691
IS-Forward KL 1 24.88 0.756
2 27.9 0.759
3 30.09 0.760
$\chi^{2}$ 1 26.7 0.745
2 28.78 0.740
3 29.97 0.739
Table 3: Response diversity of SPPO and RSPO evaluated with Self-BLEU score. The regularization temperatures of RSPO are the same as in Figure 2. RSPO methods involving reverse KL regularization have both better generation quality and diversity.
We demonstrate additional advantages introduced by regularization. We first investigate the diversity of the response cause by regularization. We first provide a motivating example with synthetic data in Appendix D.2, which shows that the unregularized self-play may converge to a collapsed response when multiple equally good responses exist. On the contrary, RSPO with maximum entropy regularization has multi-modal distribution for generation.
For LLMs, we investigate the diversity of generations by estimating the variability of the responses. We use the Self-BLEU (Zhu et al., 2018) score, where a lower score implies higher response diversity. We take the first $200$ tokens of each of the $16$ generated responses using the prompts of AlpacaEval.
The trend of Self-BLEU scores presented in Table 3 (Right) show that applying RSPO with Reverse KL increases response diversity the most, as well as the LCWRs of AlpacaEval 2.0. Although reverse KL regularization is typically associated with reduced diversity, it can, counterintuitively, enhance diversity when the high-probability region of the reference policy $\mu$ contains multiple modes—a scenario commonly arising when $\mu$ is pretrained on a diverse dataset. In such cases, the sampling-regularized optimization process with reverse KL can also induce additional modes in the learned policy distribution, thereby promoting greater diversity in responses. In contrast, IS-Forward KL yields slightly lower diversity, as its importance sampling–based implementation necessitates hard clipping for numerical stability. Compared to reverse KL, the $\chi^{2}$ divergence functions as a stronger regularizer (Huang et al., 2024), promoting diversity, albeit at a slower rate.
Finally, we assess additional aspects of response quality on the Ultrafeedback validation set using ArmoRM Wang et al. (2024b) (Table 4). Reverse KL regularization improves both truthfulness and helpfulness. Notably, while forward and reverse KL regularization individually tend to diminish instruction-following performance, their combination yields improvements across nearly all evaluated aspects, attaining the highest overall score—with particularly strong gains in Instruction Following and Truthfulness.
Methods Overall Score Instruction Following Truthfulness Honesty Helpfulness
Snorkel 0.706 0.781 0.796 0.821 0.760
SPPO 0.716 0.798 0.812 0.836 0.771
RSPO ($\chi^{2}$, $\lambda=0.1$) 0.713 0.793 0.805 0.827 0.769
RSPO (Rev. $\lambda=0.5$) 0.718 0.798 0.805 0.831 0.773
RSPO (Rev. $\lambda=1$) 0.715 0.798 0.807 0.826 0.769
RSPO (For. $\lambda=0.1$) 0.711 0.795 0.809 0.824 0.760
RSPO (For. $\lambda=0.5$) 0.713 0.793 0.815 0.826 0.749
RSPO (For.+Rev.) 0.719 0.805 0.816 0.833 0.768
Table 4: ArmoRM Evaluation. Evaluation of diverse response quality aspects using ArmoRM on Ultrafeedback validation set. The combined application of forward and reverse KL regularization leads to superior performance compared to either form of regularization applied independently.
5 Related Work
Offline RLHF with general divergence for regularization. The use of general divergence-based regularization has been explored in the context of offline alignment. $f$-DPO (Wang et al., 2023) extends Direct Preference Optimization (Rafailov et al., 2024) from reverse KL regularization to a broader class of $f$-divergences, but primarily demonstrates benefits in generation diversity. The specific effects of individual divergences—and their performance on widely-used benchmarks such as AlpacaEval—remain unexamined. $\chi$PO (Huang et al., 2024) emphasizes the theoretical importance of $\chi^{2}$ divergence for uncertainty quantification. However, the role of regularization in online iterative preference optimization, particularly its empirical impact on standard benchmarks, has yet to be studied.
Contrastive Self-Play Alignment We emphasize the distinction between our self-play approach and contrastive self-play methods including Direct Nash Optimization (DNO) (Rosset et al., 2024) and Iterative Nash Policy Optimization (INPO) (Zhang et al., 2024). These methods conduct policy optimization with a loss objective necessary but not sufficient for Mirror Descent (MD) update (Beck and Teboulle, 2003). This objective is constructed via winner-loser response comparisons similar to Direct Preference Optimization (DPO) and Identity Preference Optimization (IPO) (Azar et al., 2024). Optimizing such contrastive loss can lead to only an increase in the relative likelihood gap without necessarily enhancing the absolute probability of the preferred response (Pal et al., 2024). In contrast, our method directly approximates the MD update by converting it to an equivalent reinforcement learning problem, thereby circumventing the limitations of contrastive approaches.
6 Conclusion
In this paper, we study the regularization in self-play by proposing a framework, namely Regularized Self-Play Policy Optimization (RSPO). Based on RSPO, we can apply different regularization functions for policy updates by adding the regularization term to the loss functions, which is still guaranteed to converge to the NE of the regularized preference optimization game. In the empirical assessments, we achieve significant improvement over the base model and unregularized self-play method, SPPO. We also empirically demonstrate that regularization promotes response diversity, as well as instruction following, truthfulness, and helpfulness. These findings underscore the critical role of regularization as a fundamental component in optimizing self-play alignment.
Appendix A Proofs
In this section, we provide detailed assumptions, derivations and proofs of propositions.
Assumption A.1 (Relative Convexity of $R$ w.r.t. entropy function).
We assume the regularization function $R$ of policy $\pi$ is a $1$-strongly convex relative to entropy function. In other words, $\forall\pi,\pi^{\prime}\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$, and $\psi(\pi)=\langle\pi,\log\pi\rangle$, we have
$\langle\partial_{\pi}R(\pi)-\partial_{\pi}R(\pi^{\prime}),\pi-\pi^{\prime}\rangle\geq\langle{\partial_{\pi}\psi(\pi)-\partial_{\pi}\psi(\pi^{\prime})},\pi-\pi^{\prime}\rangle.$ (16)
Assumption A.1 constrains the class of regularization terms $R$ under which theoretical convergence guarantees can be established. Nonetheless, a broad family of divergences still satisfies this assumption, allowing RSPO to retain convergence properties in a wide range of settings. Among the divergences used in our experiments—including linear combinations—only the forward KL divergence violates this assumption. Interestingly, however, forward KL regularization is empirically observed to reduce response length. To leverage this desirable property while preserving theoretical validity, we propose a linear combination of forward and reverse KL divergences, enabling effective length-controlled generation without sacrificing convergence guarantees, and obtains the best generation quality empirically.
A.1 Proof of the Existence of regularized Nash Equilibrium
Proposition A.2.
Nash Equilibrium in the regularized game in Equation 4 exists, and it is unique.
Proof.
We prove the existence of in this section, largely following the idea of proving the existence of KL regularized Nash Equilibrium by Munos et al. [2023].
Since the utility $u(\pi,\pi^{\prime})$ is linear in $\pi$ and $\pi^{\prime}$, and the regularization function is assumed to be convex (Assumption A.1), the regularized preference is concave in $\pi$ and convex in $\pi^{\prime}$. Therefore, the existence and the uniqueness of a regularized Nash Equilibrium in Equation 4 can be directly derived from the minimax theorem [Sion, 1958]. ∎
A.2 Proof of Equivalence between MD and RSPO
Proposition A.3.
Nash-MD and Online Mirror Descent [Munos et al., 2023, Section 6] can be seen as instances of Regularized Self-Play Policy Optimization (RSPO) (Equation 6).
Proof.
In this section, we first provide derivations of how Nash-MD is equivalent to RSPO:
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}=\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu}),B=\frac{1}{2},R=D_{\text{KL}}(\pi_{\theta}\|\mu)\big{)}$ (17)
On one hand, Nash-MD practical loss [Munos et al., 2023, Section 7] is defined as
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}(\theta)$ (18)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{)}\Big{]}$ (19)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}+2\tau\log\frac{\pi_{t}(y)}{\mu(y)}\Big{)}\Big{]}$ (20)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}+\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{)}\Big{]}$ (21)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\Big{)}\Big{]}+\tau\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{]}$ (22)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\Big{)}\Big{]}+\tau\nabla_{\theta}D_{\text{KL}}(\pi_{\theta}\|\mu)$ (23)
$=$ $2\tau^{2}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\Big{(}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\frac{1}{2\tau}\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}\Big{)}\Big{)}^{2}\Big{]}+\tau\nabla_{\theta}D_{\text{KL}}(\pi_{\theta}\|\mu)$ (24)
$=$ $2\tau^{2}\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{t},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\frac{1}{2\tau}\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}\Big{)}\Big{]}^{2}+\tau\nabla_{\theta}D_{\text{KL}}(\pi_{\theta}\|\mu).$ (25)
Equation 19 is the definition of Nash-MD policy gradient. Equation 20 holds because the additional term satisfies that $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\big{[}\nabla_{\theta}\log\pi_{\theta}(y)\big{(}\log\frac{\pi_{t}(y)}{\mu(y)}\big{)}\big{]}=\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\big{[}\log\frac{\pi_{t}(y)}{\mu(y)}\big{]}=\nabla_{\theta}\log\frac{\pi_{t}(y)}{\mu(y)}=0$. Equation 23 holds due to the definition of reverse KL divergence. Equation 24 is derived by computing the integral of $\log\pi_{\theta}(y)\big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\big{)}$.
On the other hand, we show that Nash-MD and OMD can also be generalized by RSPO without external regularization, such that we can add additional regularization to existing regularized self-play methods. Nash-MD practical loss [Munos et al., 2023, Section 7] is defined as
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ y^{\prime})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{)}\Big{]}$ (26)
$=\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ y^{\prime})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\Big{)}\Big{]}$ (27)
$=\mathbb{E}_{y\sim\pi_{\theta}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi^{\mu}_{t})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\Big{)}\Big{]}$ (28)
$=\mathbb{E}_{y\sim\pi_{t}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi^{\mu}_{t})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\Big{)}\Big{]}$ (29)
$=\nabla_{\theta}\mathbb{E}_{y\sim\pi_{t}}\Big{[}\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\Big{(}\mathbb{P}(y\succ\pi^{\mu}_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}-\frac{1}{2}\Big{)}\Big{]}^{2}/2$ (30)
$=\tau^{2}\nabla_{\theta}\mathbb{E}_{y\sim\pi_{t}}\Big{[}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\frac{1}{\tau}\Big{(}\mathbb{P}(y\succ\pi^{\mu}_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}-\frac{1}{2}\Big{)}\Big{]}^{2}/2.$ (31)
Equation 26 is the definition of practical Nash-MD loss [Munos et al., 2023, Section 7]. Equation 27 holds by adding an subtracting the same element $\log\pi_{t}(y)$. Equation 28 holds due to $\mathbb{E}_{y^{\prime}\sim\pi_{t}^{\mu}}[\mathbb{P}(y\succ y^{\prime})]=\mathbb{P}(y\succ\pi_{t}^{\mu})$. The learning rate $\eta$ is originally omitted in the paper [Munos et al., 2023]. Here Nash-MD is generalized by $\mathcal{L}_{\text{RSPO}}$ with $\eta=\tfrac{1}{\tau}$ and $R=0$.
OMD is to execute $\arg\max_{\pi}\eta\mathbb{E}_{y\sim\pi}\left[\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\right]-\text{KL}(\pi,\pi_{t})$. Therefore, the loss function of the OMD update satisfies
$\nabla_{\theta}\mathcal{L}_{\text{OMD}}(\theta)$ $=-\nabla_{\theta}\eta\mathbb{E}_{y\sim\pi_{\theta}}\left[\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\right]+D_{\text{KL}}(\pi_{\theta},\pi_{t})$ (32)
$=-\nabla_{\theta}\eta\mathbb{E}_{y\sim\pi_{\theta}}\left[\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}-\log\frac{\pi_{\theta}}{\pi_{t}}\right]$ (33)
$=\eta\mathbb{E}_{y\sim\pi_{\theta}}\left[-\nabla_{\theta}\log\pi_{\theta}\Big{(}\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}-\log\frac{\pi_{\theta}}{\pi_{t}}\Big{)}\right]$ (34)
$=\frac{\eta}{2}\cdot\mathbb{E}_{y\sim\pi_{\theta}}\left[\nabla_{\theta}\Big{(}\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}-\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\Big{)}^{2}\right]$ (35)
$=\frac{\eta}{2}\cdot\mathbb{E}_{y\sim\pi_{t}}\left[\nabla_{\theta}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\Big{(}\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\Big{)}\right]^{2}.$ (36)
Equation 32 holds because the OMD update is equivalent to descending negative gradient of the feedback $\eta\mathbb{E}_{y\sim\pi}\left[\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}\right]-\text{KL}(\pi,\pi_{t})$. Equation 33 holds due to the definition of $D_{\text{KL}}$. Equation 34 holds by conducting differentiation on multiplication. The remaining equations hold due to simple algebra. Therefore, OMD can also be generalized by RSPO with $G=\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}$ and without external regularization. ∎
A.3 Proof of Proposition 3.1
Proposition 3.1.
If $R(\cdot,\mu)$ is $1$-strongly convex relative to $\psi$ (Assumption A.1), policy updated by GMMD in Equation 13 has last-iterate convergence to the following Nash Equilibrium of a regularized game:
$\max_{\pi}\min_{\pi^{\prime}}U(\pi;\pi^{\prime})-\tau R(\pi,\mu)+\tau R(\pi^{\prime},\mu).$ (37)
Proof.
According to Equation 13, GMMD is equivalent to the Algorithm 3.1 in Sokota et al. [2022]:
$z_{t+1}=\arg\min_{z\in\mathcal{Z}}\eta\left(\langle F(z_{t}),z\rangle+\alpha g(z)\right)+B_{\psi}(z;z_{t}),$ (38)
where in our setting, $z=\pi$ is the LLM policy, $F(z_{t})=-\partial_{\pi}U(\pi;\pi_{t})$ is the vector of negative partial derivatives of preference w.r.t. each component of $\pi$, $\alpha=\tau$, $g(z)$ is the regularizer $R(\pi)$, and we set $\psi(z)=z\log z$ to convert the Bregman divergence $B_{\psi}$ to KL divergence. Here $U(\pi;\pi_{t})$ is treated as a function of vector form of $\pi$, i.e., $[\pi^{0}\ \pi^{1}\ \cdots\ \pi^{|\mathcal{Y}|}]$, thus the gradient is a vector gradient where $\partial_{\pi}U(\pi;\pi_{t})=[\partial U/\partial\pi^{0}\ \ \partial U/\partial\pi^{1}\ \ \cdots\ \ \ \partial U/\partial\pi^{|\mathcal{Y}|}]$.
We then show that in our setting the following assumptions are satisfied. $F$ satisfies that for $\mu>0$ and any $z,z^{\prime}$, $\langle F(z)-F(z^{\prime}),z-z^{\prime}\rangle=0$ since $U$ is linear in $\pi$, and $F(z)-F(z^{\prime})=-\partial_{\pi}U(\pi;\pi_{t})+\partial_{\pi}U(\pi^{\prime};\pi_{t})=0$. Therefore, $F$ is Monotone and $L$-smooth. According to Assumption A.1, $g$ is $1$-strongly convex relative to $\psi$, i.e., $g(z)\geq g(z^{\prime})+\frac{g^{\prime}(z)}{\psi^{\prime}(z)}(\psi(z)-\psi(z^{\prime}))$.
Given the assumptions above, according to the Theorem 3.4. in Sokota et al. [2022], the update rule defined in Equation 38 has a last-iterate convergence guarantee to a policy $\pi^{*}$, which is the solution to the variational inequality problem $\text{VI}(\Delta^{\mathcal{X}}_{\mathcal{Y}},F+\alpha\partial g)$, i.e., $\pi^{*}$ satisfies
$\langle\partial_{\pi}\Big{(}-U(\pi;\pi^{*})+\tau R(\pi,\mu)\Big{)}\mid_{\pi=\pi^{*}},\pi-\pi^{*}\rangle$ $\geq 0,\quad\forall\pi\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$
$\Leftrightarrow\langle\partial_{\pi}\Big{(}-U(\pi;\pi^{*})+\tau R(\pi,\mu)-\tau R(\pi^{*},\mu)\Big{)}\mid_{\pi=\pi^{*}},\pi-\pi^{*}\rangle$ $\geq 0,\quad\forall\pi\in\Delta^{\mathcal{X}}_{\mathcal{Y}}.$ (39)
Equation 39 indicates that moving from $\pi^{*}$ towards any direction $\pi-\pi^{*}$ can not increase the value of the objective preference model $U(\pi;\pi^{*})-\tau R(\pi,\mu)+\tau R(\pi^{*},\mu)$ at the point of $\pi=\pi^{*}$, given the opponent is $\pi^{*}$. Therefore, by symmetry, $\pi^{*}$ is the Nash Equilibrium of the regularized preference model:
$\max_{\pi}\min_{\pi^{\prime}}U(\pi;\pi^{\prime})-\tau R(\pi,\mu)+\tau R(\pi^{\prime},\mu).$ (40)
A.4 Proof of Corollary 3.2
Proof.
We prove that RSPO in Equation 8 is equivalent to GMMD up to multiplying a constant to the gradient, leading to a regularized Nash Equilibrium. We follow SPPO to replace the samples $y\sim\pi_{\theta}$ with $y\sim\pi_{t}$ directly since they are equivalent while computing the loss before updating, and rewrite the loss equivalent to GMMD:
$\nabla_{\theta}\mathcal{L}_{\text{GMMD}}(\theta)=\nabla_{\theta}\Bigg{(}\frac{1}{2}{\mathbb{E}_{{y\sim\pi_{t}}}\Big{[}-\eta G(y;\pi_{t})+\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}{+\eta B}\Big{]}^{2}}+\tau R(\pi_{\theta},\mu)\Bigg{)}=\frac{1}{2}\nabla_{\theta}\mathcal{L}_{\text{RSPO}}(\theta).$ (41)
Therefore, according to Equation 41, RSPO is the RL implementation of GMMD, since gradients of losses are equivalent up to multiplying a constant. Then we can derive the convergence gaurantee of RSPO.
$\nabla_{\theta}\mathcal{L}_{\text{RSPO}}(\theta;G=\mathbb{P}(y\succ\pi_{t}),B=\frac{1}{2})$ (42)
$=\nabla_{\theta}\Big{(}\mathbb{E}_{y\sim\pi_{t}}\Big{[}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\eta\Big{(}\mathbb{P}(y\succ\pi_{t})-\frac{1}{2}\Big{)}\Big{]}^{2}+\lambda R(\pi_{\theta},\mu)\Big{)}$ (43)
$=\nabla_{\theta}\Big{(}\langle\pi_{t},,\big{(}-\eta\partial_{\pi}\mathbb{P}(\pi\succ\pi_{t})+\log\frac{\pi_{\theta}}{\pi_{t}}{+B}\big{)}^{2}\rangle+\lambda R(\pi_{\theta},\mu)\Big{)}$ (44)
$=2\Big{(}\nabla_{\theta}\mathbb{E}_{{y\sim\pi_{t}}}[\big{(}-\eta G(y,\pi_{t})+\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}{+B}\big{)}^{2}]\cdot\frac{1}{2}+\tau\nabla_{\theta}R(\pi_{\theta},\mu)\Big{)}$ (45)
$=2\nabla_{\theta}\mathcal{L}_{\text{GMMD}}(\theta).$ (46)
Equation 43 holds due to definition. Equation 44 holds by treating policy as a vector and rewrite the expectation in vector product form, and $\nabla_{\pi}\mathbb{P}(\pi\succ\pi_{t})\mid_{\pi=\pi_{t}}\mid_{\pi=\pi_{t}}=[\mathbb{P}(y^{0}\succ\pi_{t})\quad\mathbb{P}(y^{1}\succ\pi_{t})\quad\cdots\quad\mathbb{P}(y^{|\mathcal{Y}|}\succ\pi_{t})]^{T}$, where $y^{0},y^{1},\cdots,y^{\mathcal{Y}}$ represent all possible values of $y$. Equation 45 holds by rewriting the form of dot product as expectation. Equation 46 holds due to the equivalent loss form of GMMD in Equation 41.
Thus, according to Proposition 3.1, updating following Algorithm 1 with the above loss function has last-iterate convergence to the Nash Equilibrium of the regularized preference optimization game in Equation 4 by setting $u(\pi;\pi^{\prime})=\mathbb{P}(\pi\succ\pi^{\prime})$.
A.5 Proof of Proposition C.1
Proof.
$\pi$ is parametrized by $\theta$, $\nabla_{\theta}D_{\text{KL}}(\pi||\mu)=\mathbb{E}_{\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(y)-\log\mu(y)]^{2}/2$. This is because
$\nabla_{\theta}D_{\text{KL}}(\pi||\mu)$ $=\nabla_{\theta}\sum_{y}\pi_{\theta}(y)\cdot(\log\pi_{\theta}(y)-\log\mu(y))$ (47)
$=\sum_{y}\nabla_{\theta}\pi_{\theta}(y)\cdot(\log\pi_{{\theta}}(y)-\log\mu(y))+\sum_{y}\nabla_{\theta}\pi_{\theta}(y)$
$=\sum_{y}\pi_{{\theta}}(y)\frac{\nabla_{\theta}\pi_{\theta}(y)}{\pi_{{\theta}}(y)}\cdot(\log\pi_{{\theta}}(y)-\log\mu(y))+\nabla_{\theta}\sum_{y}\pi_{\theta}(y)$
$=\mathbb{E}_{\pi_{{\theta}}}[(\log\pi_{{\theta}}(y)-\log\mu(y))\cdot\nabla_{\theta}(\log\pi_{\theta}(y)-\log\mu(y))]$
$=\mathbb{E}_{\pi_{{\theta}}}[\nabla_{\theta}(\log\pi_{\theta}(y)-\log\mu(y))^{2}]/2.$ (48)
The first equation holds because of the definition of KL divergence. The second equation holds due to applying the product rule of differentiation. The third equation holds due to simple algebra, and the second term will then vanish because of the sum of the probabilities. The fourth equation holds because of simple algebra. ∎
A.6 Proof of Proposition C.2
Proof.
$\pi$ is parametrized by $\theta$, then $\nabla_{\theta}D_{\text{KL}}(\mu||\pi)=\mathbb{E}_{\mu}[\nabla_{\theta}\frac{\mu(y)}{\pi_{\theta}(y)}]$ because
$\nabla_{\theta}D_{\text{KL}}(\mu||\pi)$ $=\nabla_{\theta}\sum_{y}\mu(y)\cdot(\log\mu(y)-\log\pi_{\theta}(y))$ (49)
$=-\sum_{y}\mu(y)\nabla_{\theta}\log\pi_{\theta}(y)=-\sum_{y}\pi_{\theta}(y)\frac{\mu(y)}{\pi_{\theta}(y)}\nabla_{\theta}\log\pi_{\theta}(y)$
$=-\mathbb{E}_{\pi_{\theta}}\bigg{[}\frac{\mu(y)\nabla_{\theta}\log\pi_{\theta}(y)}{\pi_{\theta}(y)}\bigg{]}=-\mathbb{E}_{\pi_{\theta}}\bigg{[}\frac{\mu(y)\nabla_{\theta}\pi_{\theta}(y)}{\pi_{\theta}(y)^{2}}\bigg{]}=\mathbb{E}_{\pi_{\theta}}\bigg{[}\nabla_{\theta}\frac{\mu(y)}{\pi_{\theta}(y)}\bigg{]}.$ (50)
The first three equations hold due to the definition of forward KL divergence and simple algebra. The fourth equation comes from rewriting the forward KL following the first three equations. The fifth equation holds by taking the derivative of $\log\pi_{\theta}$. The sixth equation holds since $\frac{\nabla_{\theta}\pi_{\theta}(y)}{\pi_{\theta}(y)^{2}}=\nabla_{\theta}\frac{-1}{\pi_{\theta}(y)}$. ∎
A.7 Proof of Proposition C.3
Proof.
$\pi$ is parametrized by $\theta$, $\nabla_{\theta}D_{\chi^{2}}(\pi_{\theta}(y)||\mu(y))=\mathbb{E}_{\pi_{\theta}}\left[\frac{\nabla_{\theta}\pi_{\theta}(y)}{\mu(y)}\right]$ since
$D_{\chi^{2}}(\pi_{\theta}(y)||\mu(y))$ $=\frac{1}{2}\sum_{y}\left(\frac{\pi_{\theta}(y)}{\mu(y)}-1\right)^{2}\mu(y)=\frac{1}{2}\sum_{y}\frac{\pi_{\theta}(y)^{2}-2\pi_{\theta}(y)\mu(y)+\mu(y)^{2}}{\mu(y)}$
$=\frac{1}{2}\sum_{y}\frac{\pi_{\theta}(y)^{2}}{\mu(y)}+C(\mu)=\frac{1}{2}\mathbb{E}_{\pi_{\theta}(y)}\left[\frac{\pi_{\theta}(y)}{\mu(y)}\right]+C,$ (51)
where $C(\mu)$ is independent to $\theta$. The first two equations hold according to the definition of Chi-squared divergence. The third equation holds by separating the terms only related to $\mu$ and the term related to $\pi_{\theta}$. The fourth equation holds by rewriting the summation as the expectation. ∎
Appendix B Additional Related Work
Preference Optimization.
Large Language Models (LLMs) recently have obtained remarkable capabilities to accomplish a range of tasks [Jiang et al., 2023b, Dubey et al., 2024, DeepSeek-AI et al., 2025], generating more desirable and helpful content following the user’s intention. One of the most important methods to align LLMs with human intentions is Reinforcement Learning from Human Feedback (RLHF), maximizing a preference-based reward penalized by a reverse KL regularization term of the LLM policy and a reference model [Christiano et al., 2017, Ouyang et al., 2022, Rafailov et al., 2024, Azar et al., 2024, Xiong et al., 2024]. Since the reference model usually provides safer guidance for policy optimization [Munos et al., 2023], this regularization is crucial in RLHF to prevent over-optimization, which has been extensively studied and extended beyond KL divergence [Wang et al., 2023, Go et al., 2023, Huang et al., 2024]. In this work, we instead study the regularization problems in self-play alignment.
RLHF with General Preference Optimization (Self-Play Alignment). Azar et al. [2024] introduced the first approach for optimizing LLM policy via general preference models. Nash-MD [Munos et al., 2023] pioneered the application of self-play to general preference optimization by framing it as a two-player game. Subsequent methods have either focused on learning the NE of the original unregularized game (e.g. [Swamy et al., 2024, Wu et al., 2024, Rosset et al., 2024, Wang et al., 2024a]) or the NE of a reverse-KL-regularized preference optimization game (e.g. [Munos et al., 2023, Calandriello et al., 2024, Zhang et al., 2024]). In contrast, our work explores a broader class of divergence-based regularization techniques for self-play alignment.
Notably, our RSPO can generalize existing self-play methods. Unregularized self-play methods following the preference-based MWU can all be generalized by $\mathcal{L}_{\text{RSPO}}$ without external regularization, and thus can be regularized by simply adding regularization term to the loss functions. Based on the same exponential update rule as in SPPO, SPO [Swamy et al., 2024] is equivalent to updating policy with the loss in Equation 9. Magnetic Policy Optimization [Wang et al., 2024a], despite incorporating regularization in the policy update, periodically updates $\mu=\pi_{t}$. Consequently, it inherently follows MWU while incorporating multiple policy updates within each iteration, following
Online iterative RLHF. Iterative alignment method incorporates a reliable reward or preference model—including self-play—functions as a self-improving framework by iteratively generating new data using models and optimizing policies based on this data [Schulman et al., 2017, Ouyang et al., 2022, Bai et al., 2022, Touvron et al., 2023, Dong et al., 2024]. Moreover, extending powerful offline methods such as DPO to iterative frameworks has led to significant performance gains [Xu et al., 2023, Liu et al., 2023, Tran et al., 2023, Dong et al., 2024, Calandriello et al., 2024, Pang et al., 2024, Xiong et al., 2024, Guo et al., 2024, Tajwar et al., 2024, Cen et al., 2024, Xie et al., 2024]. In contrast, our work investigates general preference optimization through self-play from a game-theoretic perspective, shifting the objective from conventional RL optimization to the computation of NE.
Appendix C Additional Details
In this section, we provide additional details of this paper, including the algorithm descriptions of self-play alignment methods, a summarizing table for generalizing existing methods, and our implementation of regularizations.
C.1 Self-Play Alignment Algorithm
Algorithm 1 shows the overall self-play alignment process. Note that we are sampling $K$ responses per each prompt and obtain pair-wise preferences amongst them for training.
Algorithm 1 Self-Play Alignment
  Input: LLM $\pi_{\theta}$, preference model ${\mathbb{P}}$, number of iterations $T$, reference policy $\mu$, loss function for policy update conditioned on utility function $U$: $\mathcal{L}(\theta;U)$, sample size $K$.
  Initialize: $\pi_{0}=\mu$.
  for $t\in[T]$ do
     Sample prompts and responses: $x\sim\mathcal{X}$, $y_{1:K}\sim\pi_{t}$
     Get pair-wise preferences $u_{ij}={\mathbb{P}}(y_{i}\succ y_{j}),\ \forall i,j\in[K]$
     Update policy parameters $\theta=\arg\min_{\theta}\mathcal{L}(\theta;U)$, $U=[u_{ij}]\in\mathbb{R}^{K\times K}$
     $\pi_{t+1}=\pi_{\theta}$
  end for
  Output: Last-iterate policy $\pi_{T}$.
Specifically, the policy is first initialized as $\pi_{0}=\mu$. Then in each iteration $t$, the opponent is set to be the last-iterate policy $\pi_{t}$ (the reason why it is called self-play), and the responses are sampled from $\pi_{t}$ (Line 4). The pairwise preferences of the sampled responses are collected using the preference model $\mathbb{P}$ (Line 5). The policy parameters are updated by minimizing a specified loss function $\mathcal{L}(\theta;\mathbb{P})$ based on preferences over responses (Line 6). The loss function $\mathcal{L}(\theta;\mathbb{P})$ is dependent on the inherent online learning method. The main difference between these methods is the choice of loss function $\mathcal{L}(\theta;\mathbb{P})$ applied to the policy update.
C.2 Generalizing Existing Methods
Table 5 shows how the existing methods of self-play alignment can be generalized without external regularization. The algorithms introduced below share the same loss structure as in Equation 6, while their differences present in the update direction $G$, baseline $B$ and the preference model.
Loss Update Direction ($G$) Baseline ($B$) Preference Model
$\mathcal{L}_{\text{SPPO}}$ [Wu et al., 2024] $\mathbb{P}(y\succ\pi_{t})$ $0.5$ $\mathbb{P}(y\succ y^{\prime})$
$\mathcal{L}_{\text{OMD}}$ [Munos et al., 2023] $\mathbb{P}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}$ Est. $\mathbb{P}_{\tau}(y\succ y^{\prime})$
$\mathcal{L}_{\text{Nash-MD}}$ [Munos et al., 2023] $\mathbb{P}^{\mu}(y\succ\pi_{t})-\tau\log\frac{\pi_{t}(y)}{\mu(y)}$ $0.5$ $\mathbb{P}_{\tau}(y\succ y^{\prime})$
Table 5: Self-play losses $\mathcal{L}_{\text{RSPO}}$ generalizes different self-play policy optimization methods. $\mathbb{P}^{\mu}(y\succ\pi_{t})=\mathbb{P}(y\succ\pi^{\mu}_{t})$, $\pi^{\mu}_{t}$ is the geometric mixture of $\pi_{t}$ and $\mu$. We abbreviate the estimated baseline that reduce the variance of $G$ the most as est.. $\mathbb{P}_{\tau}(y\succ y^{\prime})=\mathbb{P}(y\succ y^{\prime})-\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}+\tau\log\frac{\pi^{\prime}(y^{\prime})}{\mu(y^{\prime})}$ is the regularized preference model.
C.3 Implementation of Regularization
In practice, accurately estimating the gradient of the regularizer is essential, as many commonly used divergence measures are defined as expectations over $\pi_{\theta}$. The estimation of divergences has been extensively studied and widely applied in various domains [Rubenstein et al., 2019]. For completeness, in this section, we introduce the regularization methods investigated in this study, including Reverse KL, Forward KL, and Chi-Square Divergence.
We begin by deriving the estimation of the Reverse KL divergence based on the following proposition.
Proposition C.1.
Reverse KL divergence satisfies:
$\nabla_{\theta}D_{\textit{KL}}(\pi_{\theta}||\mu)=\mathbb{E}_{y\sim\pi_{\theta}}[{\nabla_{\theta}(\log\pi_{\theta}(y)-\log\mu(y))^{2}}].$ (52)
According to Proposition C.1, we can estimate the divergence with $\mathbb{E}_{y\sim\pi_{\theta}}[{(\log\pi_{\theta}(y)-\log\mu(y))^{2}}]$.
We employ two distinct approaches to estimate the forward KL divergence. The first method utilizes importance sampling, referred to as IS-For. KL, and is derived based on the following proposition.
Proposition C.2.
The gradient of forward KL divergence satisfies that
$\nabla_{\theta}D_{\text{KL}}(\mu||\pi_{\theta})=\mathbb{E}_{y\sim\pi_{\theta}}[\nabla_{\theta}{\mu(y)}/{\pi_{\theta}(y)}].$ (53)
Therefore, we can estimate the forward KL divergence by leveraging the expectation $\mathbb{E}_{y\sim\pi_{\theta}}[{\mu(y)}/{\pi_{\theta}(y)}]$ to estimate the forward KL. Notably, to mitigate the risk of gradient explosion, we apply gradient clipping with a maximum value of $10$.
The second method for forward KL is a direct estimation of $D_{\text{KL}}(\mu||\pi_{\theta})$. To achieve this, we resample responses from the reference policy $\mu$ using the same prompts from the training dataset, constructing a reference dataset. The KL divergence is then estimated directly based on its definition by uniformly drawing samples from this reference dataset. A key advantage of this approach is that it eliminates the need for importance sampling, as each policy update iteration only requires samples from $\pi_{t}$.
Similarly, we estimate the Chi-Square divergence using $\mathbb{E}_{y\sim\pi_{\theta}}\left[{\pi_{\theta}(y)}/{\mu(y)}\right]$, based on the following proposition. Due to the presence of the ratio term, Chi-Square divergence estimation also necessitates gradient clipping to prevent instability, for which we set a clip value of $10$.
Proposition C.3.
Chi-Square divergence has gradient
$\nabla_{\theta}D_{\chi^{2}}(\pi_{\theta}||\mu)=\mathbb{E}_{y\sim\pi_{\theta}}\left[{\nabla_{\theta}\pi_{\theta}(y)}/{\mu(y)}\right].$ (54)
We also explore the linear combination of different regularization functions to leverage their complementary effects, as in offline RLHF [Huang et al., 2024]. The previously established propositions for estimating divergences can still be used in the combined regularization method.
Apart from the flexibility and simplicity of applying different regularization methods, RSPO can generalize existing self-play methods, including the unregularized ones, which enables regularizing off-the-shelf self-play methods in practice with no change on their original loss functions or hyperparameters, directly adding an external regularization term to their loss functions.
We then provide the hyperparameters of regularization temperature for each regularizer in our experiments:
Divergence Parameter(s)
Reverse KL (Rev. KL) $\lambda=0.5$
Forward K (For. KL) $\lambda=1.0$
Chi-Squared ($\chi^{2}$) $\lambda=0.1$
Importance-Sampling Forward KL (IS-For.) $\lambda=0.1$
Forward and Reverse KL (IS-For.+Rev. KL) $\lambda_{1}=0.1$, $\lambda_{2}=0.5$
Table 6: Divergences and their corresponding $\lambda$ parameters.
Appendix D Additional Experiments
In this section, we provide additional experiments, including two synthetic motivating examples and additional results on language tasks.
D.1 Regularization in Game Solving
Figure 4: Motivating Example: 20 iterations of MWU and regularized MWU with the same learning rate to solve saddle point problem $\max_{y}\min_{y^{\prime}}f(y,y^{\prime},\alpha)$, where $f(y,y^{\prime};\alpha)=\frac{\alpha}{2}{y^{\prime}}^{2}+(y^{\prime}-1)(y-1)-\frac{\alpha}{2}y^{2}$, first introduced in [Sokota et al., 2022]. We assume that we only have access to a misspecified (surrogate) preference $f(y,y^{\prime};\alpha=1)$, while the ground truth human preference is $f(y,y^{\prime};\alpha=2)$.
The regularization in the preference model is not used in all game-theoretic self-play methods. Here we investigate the necessity of regularization and offer a motivating example in Figure 4, a saddle point solving problem $\min_{x}\max_{y}\frac{\alpha}{2}x^{2}+(x-1)(y-1)-\frac{\alpha}{2}y^{2}$. There exists a reference point as the initial values of $x$ and $y$. We assume that both reference point and the Nash Equilibrium (NE) of the surrogate preference model (Surrogate NE) are close to the original NE but on different sides of the original NE.
Typically, the surrogate preference/reward models are not positively related to the reference policy. Thus, it is a reasonable abstracted example of NLHF by treating the reference point as reference policy and surrogate NE as the optimal policy obtained by optimizing the surrogate preference/reward. The results of the $20$ iterations self-play MWU with an early stopping show that regularization can be used to prevent reward over-optimization (reaching surrogate NE). A well-tuned regularization leads to faster convergence to the unknown original NE. Thus, regularization can be effective in preventing over-optimization in self-play.
D.2 Diversity on 2D Example
We offer an analysis of our method compared to unregularized self-play (SPPO) on a 2D example in Figure 5. The area with a darker color is assigned a higher reward value. We use the preference defined by the $L^{2}$ norm between two actions. We also set the reference policy to be uniform. According to the figure, the unregularized method tends to converge to a single point on the manifold of the large reward. While regularized methods have diverse sampled actions.
Figure 5: Samples in a 2D example of different iterations of SPPO (top) and RSPO (bottom) with external forward KL regularization to a uniform random reference policy. SPPO added simple external regularization that can generate multi-modal policies.
D.3 More Results on AlpacaEval-2.0 and PairRM
In Figure 6 and Table 7, we present further results of RSPO evaluated using AlpacaEval. As presented in Figure 6, mixed regularization of the forward and reverse KL resulted in the best performance, while its average response length did not exceed that of reverse KL-only regularization. When compared to various other well-known baselines including GPT-4 and Claude, RSPO-trained model initialized from Mistral-7B shows notable performance, outperforming GPT-4 0314 and Llama 3 70B Instruct in LCWR. When response lengths are ignored, our RSPO-trained 7B model even outperforms Claude 3 Opus.
Figure 6: Win rates and the average length of SPPO and RSPO with different regularization methods. From left to right, regularization methods: Reverse KL, Forward KL, Chi-Squared, Importance-Sampling Forward KL, Importance-Sampling Forward, and Reverse KL linear combination.
Regularization Temperature Iter LCWR (%) WR (%) Length
forward: 0.1 reverse: 0.5 1 23.16 21.06 1763
forward: 0.1 reverse: 0.5 2 27.91 27.38 1992
forward: 0.1 reverse: 0.5 3 35.44 38.31 2286
forward: 0.01 reverse: 0.5 1 24.63 22.57 1793
forward: 0.01 reverse: 0.5 2 28.21 28.56 2006
forward: 0.01 reverse: 0.5 3 32.24 36.77 2411
Model
AlpacaEval 2.0
LC. Win Rate Win Rate
GPT-4 Turbo 50.0 50.0
Claude 3 Opus 40.5 29.1
Mistral-7B-RSPO Iter3 35.44 38.31
GPT-4 0314 35.3 22.1
Llama 3 70B Instruct 34.4 33.2
GPT-4 0613 30.2 15.8
Mistral Medium 28.6 21.9
Mistral-7B-SPPO Iter3 28.5 31.0
Table 7: Left: AlpacaEval-2.0 performance of RSPO with different regularization temperatures. Right: AlpacaEval-2.0 performance comparison with popular models. Our model, Mistral-7B-RSPO Iter3, outperforms GPT-4 0314 and Llama 3 70B Instruct in LCWR. When only win rate is considered, our model even outperforms Claude 3 Opus.
Iter LCWR (%) WR (%) Length
1 26.36 24.04 1802
2 28.38 27.43 1909
3 29.17 29.75 2051
4 28.45 30.20 2257
5 27.93 30.11 2301
6 28.03 30.99 2435
7 25.46 28.25 2471
8 22.94 28.26 2691
9 24.47 28.57 3402
Methods RSPO (IS-For.+Rev.) Iter3 Win Rate
Snorkel (Iterative-DPO) 0.55
SPPO Iter3 0.57
SimPO 0.50
Table 8: Left: SPPO replication Iteration-wise LCWR, WR, and Length results. Overoptimization exists according to the results. Right: Pairwise win rate of RSPO on Ultrafeedback validation set rated by pairRM. RSPO has higher win rates against all the baselines.
Appendix E Others
In this section, we provide other details including compute resources, societal impacts and limitations.
E.1 Compute Resources
We conduct experiments on 8×8$\times$8 ×A100 80GB for training and single A100 80GB for evaluation.
E.2 Societal Impacts
This study introduces a novel framework for fine-tuning large language models through self-play, incorporating regularization toward a reference model. Ethical considerations may emerge if the reference model exhibits harmful behaviors, or if the preference model used for policy updates inadvertently assigns higher ratings to harmful outputs. However, drawing on prior research, we find no evidence that the proposed approach poses direct negative societal impacts.
E.3 Limitations
A theoretical limitation lies in the nature of the regularization term $R$ which is required to be relatively convex with respect to entropy (Assumption A.1). Both reverse KL divergence and $\chi^{2}$ divergence satisfy this property, whereas forward KL divergence does not. This discrepancy is evident in performance metrics such as raw win rates. Interestingly, forward KL has a beneficial side effect of reducing response length. To leverage the length reduction and reconcile the decreasing win rate, we adopt a linear combination of forward and reverse KL divergences—an approach that not only satisfies the relative convexity condition but also exploits the complementary strengths of each to achieve improved control over response length while maintaining theoretical soundness.
Empirically, our experiments are conducted using lightweight general preference models (pairRM-0.4B) and base models (Mistral-7B), which impose limitations on model expressiveness. Nevertheless, the focus of this work is on the development of a principled fine-tuning algorithm, which has already demonstrated substantial empirical improvements under these constraints. Extension to larger and more capable models is straightforward within the proposed framework.